The sum of 3 numbers is 2.
the first number minus the second plus the third is -2.
The first minus the third is 2 more than the second.
Find the numbers. I am so lost.

Answer by ewatrrr(24785)

 
Hi,
Let x, y, z represent the numbers respectively
The sum of 3 numbers is 2.
the first number minus the second plus the third is -2.
The first minus the third is 2 more than the second.
Write as we read
I. x + y + z = 2
II. x - y + z = -2
III. x - z = y + 2
I. x + y + z = 2
II. x - y + z = -2
2x + 2z = 0 0r x+=+-z
III. x - z = y + 2
-z - z = y + 2
-2z+-+2+=+y+
I. -z + (-2z-2) + z = 2
-2z - 2 = 2
-2z = 4
z = -2
x = 2, y = 2, z = -2